OpenStudy (anonymous):

Indicate the equation of the given line in standard form. The line containing the diagonal of a square whose vertices are A(-3, 3), B(3, 3), C(3, -3), and D(-3, -3). Find two equations, one for each diagonal.

OpenStudy (e.mccormick):

The slope of a line is the change in y over the change in x. Or:\[\frac{y_1-y_2}{x_1-x_2}If you were to take opposite corners of the square, they would be the pints you used to do this.Seeing this square, I know some other ways to find the slope of the diagonals, but once you do have the slope, all you need to do is pick any point and use it to find the equation of the line. If you have a known \((x,y)\) in the form of \((h,k)\) and have solved for the slope of m, then the line can be found by using \(y-k=m(x-h)\). In your case, the \((h,k)\) would be one of the 4 corners of the square. Just make sure you pick one on the line you are solving for.